OpenStudy (anonymous):

During the period 1985–2012, the projected enrollment B (in thousands of students) in public schools and the projected enrollment R (in thousands of students) in private schools can be modeled by B = –18.53t2 + 975.8t + 48,140 and R = 80.8t + 8049, where t is the number of years since 1985.

OpenStudy (anonymous):

Hold on . I wrote this wrong . One sec.

OpenStudy (anonymous):

During the period 1985–2012, the projected enrollment B (in thousands of students) in public schools and the projected enrollment R (in thousands of students) in private schools can be modeled by B = –18.53t^2 + 975.8t + 48,140 and R = 80.8t + 8049, where t is the number of years since 1985.

OpenStudy (anonymous):

Write an equation that models the difference in the projected enrollments for public schools and private schools as a function of the number of years since 1985.

OpenStudy (anonymous):

Find the difference in projected enrollments for public schools and private schools in 2005.

OpenStudy (paxpolaris):

\[B = –18.53t^2 + 975.8t + 48,140\]\[ R = 80.8t + 8049\] \[B-R=–18.53t^2 + 975.8t + 48,140 - \left( 80.8t + 8049 \right)\]

OpenStudy (anonymous):

Thank you.